# Mission-Aligned Learning-Informed Control of Autonomous Systems: Formulation and Foundations

### v1

@@ 7 Conclusion
The increasing interest of deploying AI in real world scenarios, from autonomous vehicles and drones to robotics in manufacturing and care, presents the necessity of enforcing reliability through first principles domain knowledge. Increasingly, there is a complex layered taxonomy of tasks and behaviors, and interpretability is essential for fostering trust and robust performance. This paper presents a first attempt at simultaneously integrating classical planning, which encodes crisp goals in a solid logical framework, through to MPC, the most state of the art the technology for fast real time control, while incorporating classic RL tools for learning the rewards from the environment. We defined the entire end-to-end Scheduling to Planning to Control, including both the operation and the co-learning across layers. We hope the work opens up a productive research program towards the management and control of autonomous agents.

### v2

@@ 10 Conclusion
Summary
This paper addressed a fundamental tension in deploying learning-enabled robots in safety-critical environments: black-box reinforcement learning can discover effective policies but offers no guarantees on safety, interpretability, or predictability — properties that are non-negotiable in domains such as hospital care. We presented a multilevel control architecture that achieves the adaptive personalisation of RL while maintaining the deterministic, interpretable structure required for autonomous systems operating alongside vulnerable populations.
The bilevel structure couples an outer loop that performs classical planning - an interpretable and reliable algorithmic toolkit for choosing discrete combinatorial action sequences that achieve a goal. Each action defines a parametrization of an optimal control problem. An MPC controller defines continuous actuator operation of an autonomous system that appropriately incorporates known physics engineering. Learning is performed between the layers through real-time reward signals and discrete-continuous fuzzy map discrepancies. These propagate, through MPC sensitivities, as gradient information, closing the entire system loop.
We validated the system in a hospital service robot domain with two structurally distinct task types - medication delivery and meal preparation - across five patient profiles spanning different regions of a 5-dimensional preference space. Consistently successful task completion and adaptation to patient preferences was observed in the simulation validation.
Broad Conclusions
The increasing interest of deploying AI in real world scenarios, from autonomous vehicles and drones to robotics in manufacturing and care, presents the necessity of enforcing reliability through first principles domain knowledge. Increasingly, there is a complex layered taxonomy of tasks and behaviors, and interpretability is essential for fostering trust and robust performance. The current default of scaling Deep RL for operating these autonomous systems presents significant risks to safety as well as exhibiting limitations in agreeable deployment in human-sensitive settings due to its opaque nature. In addition, the requirement of increasingly deeper models and training computation presents unsustainable energy and economic costs, as far as the desired scale of deployment of such systems. Incorporation of crisp interpretable planning together with physics-driven control is, in the authors’ view, essential towards comprehensive reliability and satisfying performance of autonomous AI systems.
This paper presents a first attempt at simultaneously integrating classical planning, which encodes crisp goals in a solid logical framework, through to MPC, the most state of the art the technology for fast real time control, while incorporating classic RL tools for learning the rewards from the environment. We defined the entire end-to-end Scheduling to Planning to Control, including both the operation and the co-learning across layers. We hope the work opens up a productive research program towards the management and control of autonomous agents.
The potential for future work extending the presented framework is vast - mathematical theory proving guarantees of approximate optimality, stability, sample complexity and other criteria, validation in physical settings, synergistic bilevel algorithmic development, expansion across use cases. We hope that this work triggers a powerful and effective research and development program towards reliable and effective autonomous systems broadly.